Problem 1 (25 points) You are a consultant that has been asked by a law firm to compare their traditional costing system with an activity based costing system. You have gathered the following information relative to the operations of the firm INFORMATION RELATED TO OPERATIONS OF LAW FIRM Partner Labor Associate Labor Number of attorneys 20 5 Annual billable hours per attorney 2,000 2,000 100,000 Annual compensation per attorney 250,000 S S The annual estimated overhead per activity center is provided below. INFORMATION RELATED TO ESTIMATED OVERHEAD TO EACH ACTIVITY CENTER Secretarial Support Legal Support 2.500,000 S Annual estimated overhead 1,000,000 You are going to analyze two jobs for the firm: the Smith job, which will require 50 partner hours and 20 associate hours; and the Jones job, which will require 15 partner hours and 55 associate hours
REQUIRED: ( Using a traditional costing system, determine the following amounts Cost of each job, assuming that overhead is allocated to jobs using (a) total professional labor hours as the activity base. Round your calculations to the nearest whole dollar Amount that should be billed to each customer, assuming that the firm bills customers at a 25% markup on cost. Round your calculations to the nearest whole dollar (b) Using an activity based costing system, determine the following amounts Cost of each job, ass uming that overhead is allocated to jobs by using one rate for legal support using associate labor hours as the activity base and another rate for secretarial support using partner labor hours as the activity base. Round your calculations to the nearest whole dollar Amount that should be billed to each customer, assuming that the firm bills customers at a 25% markup on cost. Round your calculations to the nearest whole dollar. (2) (a) (b)
Answer #1
1. Computation of Job cost and billed amount for Smith Job and Jones Job Jobs using company's traditional costing system:-
Computation of overhead rate :-
Estimated Total Overheads 2500000+1000000 = $                3,500,000
Total professional labor hours (2000*5)+(2000*20) = 50000
Overhead rate = Estimated Total Overheads
Total professional labor hours
= $3,500,000
50000
= $70.00 per professional labor hour
(a) Computation of Cost :-
Job Activity Driver Plantwide OH Rate Overhead Cost
a b c d = b×c
Smith Job 70 Professional hours $70.00 per professional labor hour $4,900
Jones Job 70 Professional hours $70.00 per professional labor hour $4,900
(b) Computation of Amount that should be billed to customer :-
Job
Smith Job Jones Job
Cost $                        4,900 $                                   4,900
Add: Markup @25% $                        1,225 $                                   1,225
Billed amount $                       6,125 $                                  6,125
2. Computation of Job cost and billed amount for Smith Job and Jones Job Jobs using activity based costing system:-
Computation of Cost per driver :-
Activity cost pool Activity measures (Cost Drivers) Estimated Overhead Cost Expected Activity Cost per driver
No. of attorneys Annual billable hours per attorney Total
a b c d e f = d*e g = c/f
Legal Support Associate labor hours $                  2,500,000                        20                            2,000                                     40,000 $                                 62.50
Secretarial Support Partner labor hours $                  1,000,000                          5                            2,000                                     10,000 $                                    100
Total Overhead Cost $                  3,500,000
(a) Computation of Cost :-
Overhead Assigned Activity Driver Cost per driver Total Overhead Cost
a b c d = b×c
Smith Job
Legal Support                                    20 Associate labor hours $                                  62.50 $                                 1,250
Secretarial Support                                    50 Partner labor hours $                                100.00 $                                 5,000
Total Cost $                                 6,250
Jones Job
Legal Support                                    55 Associate labor hours $                                  62.50 $                                 3,438
Secretarial Support                                    15 Partner labor hours $                                      100 $                                 1,500
Total Cost $                                 4,938
(b) Computation of Amount that should be billed to customer :-
Secretarial Support
Jones Job Legal Support
Cost $                        6,250 $                                   4,938
Add: Markup @25% $                        1,563 $                                   1,234
Billed amount $                       7,813 $                                  6,172

Feel free to ask any clarification, if required. Please provide feedback by thumbs up, if satisfied. It will be highly appreciated. Thank you.